Find the Vr of the vectors given below using both the head-to-tail and component methods.

 

An athlete runs 65.0 km SE the proceeds 50.0 km NE before running 45.0 km NW and 60.0 km W.

Scale = 1.0 cm:10 km


1
Expert's answer
2021-10-10T16:02:48-0400

By the component method:


"V_{1S}=65\cos45°=46\text{ km},\\
V_{1E}=65\sin45°=46\text{ km}.\\
V_{2S}=-50\cos45°=-35\text{ km},\\
V_{2E}=50\sin45°=35\text{ km}.\\
V_{3S}=-45\cos45°=-32\text{ km},\\
V_{3E}=-45\sin45°=-32\text{ km}.\\
V_{4S}=60\cos90°=0\text{ km},\\
V_{4E}=-60\sin90°=-60\text{ km}.\\\space\\
V_{rS}=V_{1S}+V_{2S}+V_{3S}+V_{4S}=-21\text{ km},\\
V_{rE}=V_{1E}+V_{2E}+V_{3E}+V_{4E}=-11\text{ km},\\\space\\
V_r=\sqrt{V^2_{rE}+V^2_{rS}}=24\text{ km}.\\\space\\
\theta=\arctan\frac{21}{11}=62°\text{ North of West.}"